State clearly by what is meant by two events being statistically independent. b. In 2015, a company estimated that 65% of its customers were female and, in that year, three customers entered one of their shops. Assuming that these customers were not related or acquaintances, find the probability that they were: i. all female ii. all male iii. at least one male c. In a certain factory which employs 10,000 men, 1% of all employees have a minor accident in a given year. Of these, 40% had safety instructions whereas 90% of all employees had no instructions. What is the probability of an employee being accident free: i. Given that he had no safety instructions ii. Given that he had safety instructions
1
Expert's answer
2020-03-26T14:35:11-0400

a. Let "<\Omega,F,P>" is a probability space. Random events "A" and "B" from "F" are said to be independent if "P(A\bigcap B)=P(A)\cdot P(B)."

b. i) Let random event "A" is "the first customer is female", random event "B" is "the second customer is female", random event "C" is "the third customer is female". Then "P(A\bigcap B\bigcap C)=P(A)\cdot P(B)\cdot P(C)=(0.65)^3\approx 0.275"

(events "A, B \text{ and } C \text{ are collectively independent)}."

ii) Probability that the customer is male equals "1-0.65=0.35."

Let random event "A" is "the first customer is male", random event "B" is "the second customer is male", random event "C" is "the third customer is male". Then "P(A\bigcap B\bigcap C)=P(A)\cdot P(B)\cdot P(C)=(0.35)^3\approx 0.043"

(events "A", "B" and "C" are collectively independent).

iii) We have to find probability that at least one (the first, the second or the the third) customer is male:

"P=1-0.275=0.725" (events "all three customers are female" and "at least one customer is male" are opposite).

c. Random events:

"A" --- "did not have an accident"

"B" --- "had an accident"

"C" --- "had instructions"

"D" --- "did not have instructions"

We have:

"P(A)=0.99\\
P(B)=0.01\\
P(C)=0.1\\
P(D)=0.9\\
P(C|B)=0.4\\
P(D|B)=0.6"

We should find "P(A|D)" and "P(A|C)."

"P(A|D)=\frac{P(A)\cdot P(D|A)}{P(D)}\\
P(D)=P(A)P(D|A)+P(B)P(D|B)\\
0.9=0.99P(D|A)+(0.01)(0.6)\\
P(D|A)\approx 0.903.\\
i) P(A|D)=\frac{(0.99)(0.903)}{0.9}=0.9933.\\
P(A|C)=\frac{P(A)\cdot P(C|A)}{P(C)}\\
P(C|A)=1-P(D|A)=1-0.903=0.097\\
ii) P(A|C)=\frac{(0.99)(0.097)}{0.1}=0.9603."